Calculate delta h delta s delta g and t of the reaction s+o2=so2
1
Expert's answer
2019-12-23T05:05:50-0500

"\u2206H=H(so2)- H (s+o2)"

∆H=296.81-(36.36+48) =212.45 J/mol

"\u2206S= S(products)- S(reactants)"

∆S= S(so2) -S(s+o2)=248.21-( 32.054+205.0) =11.156 J/K.mol

"\u2206G=\u2206H- T*\u2206S"

∆G=212.45- 273.15*11.156 = -2834.81J/mol